OpenStudy (anonymous):

What is the standard deviation of the data set given below? 4, 7, 8, 9, 12 A. 6.8 B. Rad 6.8 C. Rad 8.5 D. Rad 2.5 E. 2.5 F. 8.5

OpenStudy (anonymous):

@lizz123

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please youhave to find the mean value first

OpenStudy (lizz123):

add 3+7+8+9+12

OpenStudy (lizz123):

then divide it by 5

OpenStudy (michele_laino):

@apexhelpp try please!

OpenStudy (anonymous):

22.2

OpenStudy (lizz123):

nope. 3+7=10+8=18+9=27+12=39 39/5=7.8

OpenStudy (anonymous):

wait that isnt the answer right?

OpenStudy (lizz123):

ik

OpenStudy (anonymous):

then what do we do

OpenStudy (lizz123):

@Michele_Laino

OpenStudy (anonymous):

my answer was 8.414 so would 8.5 be the answer?

OpenStudy (michele_laino):

Sorry I got mean=8